OpenStudy (anonymous):

Brenda purchased a total of five computers for $3,300. She received a 20% trade discount. What did Brenda pay for the computers?

jimthompson5910 (jim_thompson5910):

20% of $3,300 = (20/100) * 3300 20% of $3,300 = (0.20) * 3300 20% of $3,300 = 660

jimthompson5910 (jim_thompson5910):

So we save $660 which is the amount we take off of the initial price 3300 - 660 = 2640 So she paid $2,640

OpenStudy (anonymous):

Thanks for the help!

jimthompson5910 (jim_thompson5910):

you're welcome